# Bridging Online Behavior and Clinical Insight: A Longitudinal LLM-based Study of Suicidality on YouTube Reveals Novel Digital Markers

### v1

@@ A.5 Dateset statistics of pre- and post event
Control (Major Life Event) Traumatic Event 43 192 88 330 119 50 132
e.g., death or suicide attempt of
parent/child/sibling/close relative
Control (Personal) Middle of Uploads 95 78 42 99 76 42 97
midpoint of upload dates was determined
by sorting the unique dates and selecting
the median as the reference point
Table A.2: Descriptive statistics of YouTubers’ number of uploads pre- and post reference events across four groups - Attempted (During), Attempted (Before), Control (Major Life Events), Control (Personal). The statistics show the mean, median, and standard deviation of uploads in pre- and post-event periods. Comments provide details about event types and reference point calculations.
Interval Description Mean (days) Median (days) SD Min Max
Last upload before $\rightarrow$ First upload after 131.74 59.0 161.48 5 702.0
Last upload before $\rightarrow$ Attempt 83.02 20.5 160.84 1 953.0
Attempt $\rightarrow$ First upload after 72.42 20.5 129.74 1 669
Table A.3: Summary of Temporal Intervals Around Suicide Attempts in the Attempted (During) Group
Interval Description Mean (days) Median (days) SD Min Max
Last upload before $\rightarrow$ First upload after 214.26 39.0 565.95 2 3255
Last upload before $\rightarrow$ Major Life Event 129.77 6.0 506.89 1 3187
Major Life Event $\rightarrow$ First upload after 84.49 17.0 172.63 1 903
Table A.4: Summary of Temporal Intervals Around Traumatic Events for the Control (Major Life Event) Group
Appendix B Appendix B - Procedures
B.1 Topic modeling
B.1.1 Technical Pipeline
First, we obtained embeddings of textual segments using Sentence-BERT (SBERT), a modification of the Bidirectional Encoder Representations from Transformers (BERT), network to efficiently create short text embeddings. Uniform Manifold Approximation Projection (UMAP) [31], was used for dimensionality reduction, reducing the embeddings from $300$ to $5$ dimensions. Next, the embeddings were clustered with the Hierarchical Density-based spatial clustering of applications with noise (HDBSCAN) algorithm [30]. HDBSCAN identifies and excludes noisy data points from the clusters, contributing to greater clarity of the topic model. In addition, HDBSCAN provides probability distribution over the topics, as a result of soft clustering (mimicking traditional topic modeling techniques like Latent Dirichlet Allocation (LDA) [5]), in which each document is modeled as a distribution of topics. Finally, each cluster is represented by its top $10$ words and $3$ most representative textual segments, which were used as input to an LLM (specifically LLama-7B [48]), and facilitated in naming the clusters (see B.1.2 for full prompt used in naming the clusters).
While training different topic models, we explored various hyper-parameters and concluded with the following set of parameters:
Table B.1: Model and General Hyper-parameters
Hyperparameter Value
Model meta-llama/Llama-2-7b-chat-hf
Minimum Document Frequency 0.001
Number of Topics auto
N-gram Range (1, 2)
Minimum Topic Size 180
Table B.2: UMAP Hyper-parameters
Hyperparameter Value
Metric cosine
Minimum Distance 0
Low Memory true
Number of Neighbors 15
Number of Components 5
Random State 33
Table B.3: HDBSCAN Hyper-parameters
Hyperparameter Value
Metric euclidean
Minimum Samples 60
Prediction Data true
Minimum Cluster Size 180
Generate Minimum Spanning Tree true
Cluster Selection Method eom
B.1.2 Prompts for Topic Modeling and Matching
This section contains the prompt we used as part of the topic modeling and matching processes. The first prompt used in the LLM-based topic modeling pipeline is described in B.1.2. The LLM (LLama-7B) was provided with the 10 top TF-IDF terms and 3 most representative textual segments, and tasked with generating concise topic labels based on the provided content. The second prompt used in the matching process, as part of the LLM-refined matching by GPT-4 (§A.4.4) and can be seen in B.1.2.
Topic Modeling Naming - <s>[INST] <<SYS>> You are a helpful, respectful and honest assistant for labeling topics. <</SYS>> I have a topic that contains the following documents: - Traditional diets in most cultures were primarily plant-based with a little meat on top, but with the rise of industrial-style meat production and factory farming, meat has become a staple food. - Meat, but especially beef, is the word food in terms of emissions. - Eating meat doesn’t make you a bad person, not eating meat doesn’t make you a good one. The topic is described by the following keywords: ’meat, beef, eat, eating, emissions, steak, food, health, processed, chicken’. Based on the information about the topic above, please create a short label of this topic. Make sure you only return the label and nothing more. [/INST] Environmental impacts of eating meat [INST] I have a topic that contains the following documents: [DOCUMENTS] The topic is described by the following keywords: ’[KEYWORDS]’. Based on the information about the topic above, please create a short label of this topic. Make sure you only return the label and nothing more. [/INST]
B.2 Temporal Alignment
Figure B.2: Overview of the Temporal Analysis Pipeline.
In Figures 9 and 10, we analyzed the temporal dynamics of the topics YouTube Engagement and Mental Health Struggles pre- and post-event. As part of this analysis, we encountered the challenge of aligning different users’ uploads to a mutual timeline. This alignment was performed relative to a specific reference point, which varied by group. For the Attempted (During) and Control (Major Life Event) groups, we used the manually extracted attempt or major life event date, while for the Attempted (Before) and Control (Matches) group, we used the midpoint of upload period. As part of this analysis, we excluded uploads from the attempt/major life event date for all applicable users.
B.3 Temporal Testing Procedure
Procedure 1: Within-Group Comparison (Left-Side Figures)
Procedure 2: Between-Group Comparison (Right-Side Figures)
Appendix C Appendix C - Results
C.1 Topic Modeling of YouTube Channels
Our analysis leveraged an LLM-based topic modeling approach to analyze the transcripts from Attempted (Before) and Attempted (During) channels. This analysis resulted in the clustering of these transcripts and the identification of 166 topics, which were automatically named using the Llama-2-7B model [48] (refer to B.1.2 for the full prompt used). The spatial distribution of the topics’ embedding vectors is illustrated in Figure C.1.
Our quantitative analysis of topic quality found that topics with similar names, such as Mental Health Struggles and Mental Health Awareness, were semantically related. This relationship is also reflected in their high cosine similarity [44], with a score of 0.72, indicating that topics with similar LLM-generated names also tend to have similar underlying content.
Additionally, certain topics were user-specific, appearing in less than 10% of the users (e.g., Filipino Culture and Language), while others were more prevalent, appearing in more than 70% of the users’ videos (e.g., YouTube Engagement). A detailed description of the topics, including their names, frequencies, most representative textual segments, and correlations between topics, is provided in the Supplementary Materials Appendix D.
C.2 Topic Modeling of Narrative Suicide Stories
Figure C.2: Attempt Stories Inter-topic Distance Map showing the results of topic modeling trained on the attempt stories. The map visualizes clusters of different topics, named by LLM, where the size of each topic bubble indicates the number of document chunks classified under that topic. The figure also includes zoomed-in views of a cluster of topics centered around mental health experience -(e.g., Mental Health Support, Seeking Help and Support, Religious Mourning and Comfort)
The primary topic model analyzed in this study (§6) was trained on data from the treatment group, including all uploads from both the Attempted (Before) and Attempted (During) groups (including the attempt narrative videos).
C.3 Supplementary Statistical Analysis
C.3.1 Bottom-up Approach
Figure C.3 illustrates the word-cloud of the 5 topics identified in this analysis, and Figure C.4 illustrates the average dynamics of the topics Fashion and Style, LGBTQ+ Identity and Acceptance, and Rape and Sexual Assault.
Figure C.3: Word clouds based on the TF-IDF scores of the words in the five topics identified as predictive of the period prior to a suicide attempt (§3.1.1). The size of the words in each cloud is proportional to their TF-IDF score.
Figure C.4 presents the mean aggregated values before and after the reference event, which reveals that the Fashion and Style topic distinguishes the two Attempted groups, showing lower prevalence than the Controls. A review of the demographic information (Table 1) indicates that the proportion of females in the groups does not differentiate between the Attempted and Control groups, suggesting that this factor does not account for the observed difference.
Figure C.4: Fashion and Style (left), LGBTQ+ Identity and Acceptance (middle), and Rape and Sexual Assault (right): Average Comparison Across Groups. The figures present the average topic values across four groups comparing Pre- and Post-Event (suicide attempt, major life event, or synthetic condition). These results highlight the observed trends of lower engagement with Fashion and Style topics in the Attempted groups and higher engagement with LGBTQ+ Identity and Acceptance topics in the Attempted (Before) group. These topics were identified in answering §3.1.1.
C.3.2 Effects Analysis
As part of both the bottom-up and hybrid approaches, we performed a two-way mixed-effects analysis of variance (ANOVA), with time (pre- vs. post-reference point) as a within-subjects factor and group as a between-subjects factor. The four groups included: Attempted (Before), Attempted (During), Control (Major Life Event), and Control (Matches). The goal was to identify topics that differed across groups, changed over time, or showed differential changes across groups, while the main findings are presented in the (§3.1.2) and (§3.4.2).
The results of the two-way mixed-effects ANOVA are summarized in Table C.1, which presents the significant topics identified through the bottom-up (data-driven) approach, and Table C.4, which reflects the outcomes of the hybrid approach that incorporates expert clinical input.
To further investigate these findings, we conducted post-hoc Tukey HSD comparisons for both main and interaction effects:
Group-level effects: Tables C.2 and C.5 report pairwise comparisons between groups, averaged across time. These tests assess whether certain groups consistently exhibited higher or lower topic prevalence, irrespective of temporal dynamics.
Interaction effects: Tables C.3 and C.6 present comparisons of temporal change (i.e., the difference between pre- and post-event scores) across groups. These analyses test whether the direction or magnitude of topic change over time significantly differed between group pairs.
Overall temporal changes: Table C.7 provides the mean topic probabilities for each condition (pre and post) aggregated across all users, offering a descriptive overview of how topic prevalence shifted over time regardless of group membership.
Topics marked with *** were identified by both bottom-up and hybrid approaches.
Table C.1: Two-Way ANOVA summary (Bottom-up-identified topics)
Topic Effect df1 df2 Mean Square F Statistic p-value
Mental Health Struggles *** group 3 311 0.16 4.25 <0.01
condition 1 311 0.11 18.45 <0.01
Interaction 3 311 0.02 3.47 0.02
YouTube Engagement group 3 311 0.59 10.55 <0.01
condition 1 311 0.10 10.18 <0.01
Interaction 3 311 0.05 5.14 <0.01
LGBTQ+ Identity and Acceptance *** group 3 311 0.01 0.28 0.04
Rape and Sexual Assault *** Interaction 3 311 0.001 2.68 0.047
Fashion and Style group 3 311 1.4 27.8 <0.01
Table C.2: Tukey HSD post-hoc Group Comparisons (Bottom-up-identified topics).
Group difference = Group2 mean - Group1 mean.
Topic Group 1 Group 2 Group Difference 95% CI p-adj
Mental Health Struggles *** Attempted (Before) Control (Matches) -0.06 [ -0.11, -0.01 ] 0.01
Attempted (Before) Control (Major Life Event) -0.07 [ -0.13, -0.01 ] 0.03
YouTube Engagement Attempted (Before) Control (Matches) 0.10 [ 0.04, 0.16 ] <0.01
Attempted (Before) Control (Major Life Event) 0.11 [ 0.03, 0.18 ] <0.01
Attempted (During) Control (Matches) 0.12 [ 0.04, 0.2 ] <0.01
Attempted (During) Control (Major Life Event) 0.13 [ 0.03, 0.22 ] <0.01
Fashion and Style Attempted (Before) Control (Matches) 0.18 [0.12, 0.23] <0.01
Attempted (Before) Control (Major Life Event) 0.14 [0.06, 0.21] <0.01
Attempted (During) Control (Matches) 0.16 [0.08, 0.24] <0.01
Attempted (During) Control (Major Life Event) 0.12 [0.03, 0.21] <0.01
Table C.3: Tukey HSD post-hoc Group × Time interaction (Bottom-up-identified topics).
Group and Time Difference = (Group2 post - pre) - (Group1 post - pre).
Topic Group 1 Group 2 Group and Time Difference 95% CI p-adj
Mental Health Struggles *** Attempted (During) Attempted (Before) -0.06 [-0.12, -0.08] <0.01
YouTube Engagement Attempted (During) Attempted (Before) -0.08 [-0.14, -0.01] 0.02
Attempted (During) Control (Matches) -0.11 [-0.18, -0.04] <0.01
Attempted (During) Control (Major Life Event) -0.08 [-0.17, -0.01] 0.04
Table C.4: Two-Way ANOVA summary (Hybrid-identified topics)
Topic Effect df1 df2 Mean Square F Statistic p-value
Autism and Disability condition 1 311 0.004 6.23 0.01
Underage Smoking and Drug Use condition 1 311 0.006 13.63 <0.001
Homelessness condition 1 311 0.040 17.87 <0.001
Interaction 3 311 0.009 3.95 0.009
Anxiety and Panic Attacks group 3 311 0.008 4.28 0.006
Interaction 3 311 0.002 3.45 0.017
Social Media Addiction group 3 311 0.008 3.72 0.012
condition 1 311 0.011 8.86 0.003
Addiction Recovery condition 1 311 0.030 19.54 <0.001
Sleep Disorders and Habits group 3 311 0.007 3.42 0.018
condition 1 311 0.011 10.12 0.002
Trauma-Induced Mental Health Disorders condition 1 311 0.009 9.43 0.002
Medication Use and Addiction group 3 311 0.016 3.66 0.013
condition 1 311 0.007 10.47 0.001
Table C.5: Hybrid Approach: Tukey HSD Multiple Comparison of Groups
Topic Group 1 Group 2 Group Difference 95% CI p-adj
Anxiety and Panic Attacks Attempted (Before) Control (Matches) -0.01 [ -0.02, -0.003 ] 0.01
Medication Use and Addiction Attempted (Before) Control (Matches) -0.02 [ -0.03, -0.002 ] 0.03
Table C.6: Tukey HSD post-hoc Group Comparisons (Hybrid-identified topics).
Group Difference = Group2 mean - Group1 mean.
Topic Group 1 Group 2 Group and Time Difference 95% CI p-adj
Homelessness Attempted (Before) Control (Matches) 0.02 [0.001, 0.05] 0.04
Attempted (Before) Control (Major Life Event) 0.03 [0.004, 0.06] 0.02
Anxiety and Panic Attacks Attempted (Before) Control (Major Life Event) 0.02 [0.01, 0.04] 0.01
Table C.7: Mean topic probabilities over all users pre- and post-event (Bottom-up and Hybrid)
Topic Mean (Pre) Mean (Post)
Medication Use and Addiction 0.03 0.03
Homelessness 0.05 0.07
Trauma-Induced Mental Health Disorders 0.03 0.04
Fashion and Style 0.16 0.19
Addiction Recovery 0.05 0.07
Underage Smoking and Drug Use 0.01 0.02
Sleep Disorders and Habits 0.04 0.05
YouTube Engagement 0.41 0.44
Autism and Disability 0.03 0.03
Mental Health Struggles 0.17 0.20
C.3.3 Control Measures Analysis
Figure C.5: Temporal trends around the start of the COVID-19 pandemic in the USA (March 1, 2020), and around Reference Event (suicide attempt, major life event and midpoint of upload) in the topics COVID-19 Pandemic (top) and Mental Health Struggles (bottom).
To corroborate that the observed topic shift was specifically associated with suicide attempts and not driven by other external factors, we examined topic dynamics around the COVID-19 pandemic, which served as an absolute reference point unrelated to individual users. We performed the same between analyses as in 3.2.
This analysis revealed an expected pattern as seen in Figure C.5: both the treatment and control groups exhibited an sudden increase in the topic COVID-19 Pandemic around the pandemic’s onset (March 1st 2020), that didn’t show any significant difference between the groups, and a gradual increase in the topic Mental Health Struggles, that is different in the temporal dynamics than the pattern detected by the suicide attempt.
Appendix D Appendix D - Topic Information
This section includes two tables with prevalent topics identified in this study. Table LABEL:tab:topic_info describes the topics identified in our bottom-up (§3.1) and hybrid approaches (§3.4). Table D.2 presents topics that are highly correlated and have similar semantic meaning.
Table D.1: Overview of topic information, including: Count - the number of transcript segments assigned to each topic, Llama2 Name - the name generated by the Llama-7b model (see Section B.1.2), Representative Transcript Segment - the transcript segment closest to the topic’s center and most representative of it, and Top Words - the 10 words with the highest TF-IDF scores. Red topic names indicate topics that were identified in our bottom-up approach. Blue topics were found in our hybrid approach, and brown indicates topics that were found by both approaches. Corresponding * indicates significance with regards to temporal changes in Attempted (During).
@@ Appendix D Appendix D - Topic Information
1041 I’m going to pop in something really sour into my mouth, or I’m going to smell something that’s going to get my brain out of here. Any tips on, uh, anxiety panic attacks, panic attacks are hard to ground yourself from and they’re hard to stop yourself. There is huge differences between anxiety attacks and panic attacks. One of the biggest ones is there’s no trigger necessarily behind a panic attack. Anxiety and Panic Attacks * anxiety,panic,attacks,panic attacks,panic attack,attack,anxiety attack,anxious,anxiety attacks,anxiety like
@@
3363 So Amanda, how are you today? I’m doing okay. So I can tell you a little bit about Obsessive-Compulsive Disorder, or OCD. It’s a mental illness and it’s one of the anxiety disorders. And so there’s two parts. There’s the obsessions and there’s the compulsions. So the obsessions, those are repetitive and intrusive thoughts. It can be urges or images that cause a lot of anxiety in an individual. And then the second part are the compulsions. Obsessive-Compulsive Disorder (OCD) ocd, compulsions, intrusive, intrusive thoughts, people ocd, ocd like, obsessions, ocd just, obsessive compulsive, obsessive
1036 Hi guys welcome to educating Shani I’m Shani and I’m recovering from an eating disorder. Hi I’m Michelle. Wait what am I supposed to say? Say hi Shani. Hi Shani. Hi! Welcome to… Start over. Hi guys welcome to educating Shani I’m Shani and I’m recovering from an eating disorder. I’ve been in this chair for like two. Yeah well it won’t be two hours after editing. Eating Disorder Recovery shani,shanny,hi shani,welcome educating,educating,disorder hi,educating shani,shani shani,shani hi,shani recovering
219 She becomes a rebellious sure goes off to college and Cal basically just goes to high school and does his own thing Amore stays in it or Sonny stays in the crib and that’s it right Basil It’s so weird cuz Basil I feel like Basil Basil and Amore saw Mary kill herself or like walked in on Mary Which is why it’s like they’re both kind of like I would say traumatized a little bit Loss and Trauma in a Small Community basil,aubrey, mary, sonny, amore, kel, mary died, photo album, sunny, hero, basil basil, mori,right aubrey, aubrey aubrey, hector
@@
Table D.2: Correlation Matrix of 30 Topics Based on Cosine Similarity: The matrix, computed using cosine similarities of the topic embeddings, demonstrates that topics with high correlation are also semantically similar, as anticipated.

### v2

@@ D.2.4 Activity Analysis (Upload Frequency)
Interval Description Attempted (During) Group Control (Major Life Event) Group
Mean Median SD Min Max Mean Median SD Min Max
Last upload before $\rightarrow$ First upload after 131.74 59.0 161.48 5 702 214.26 39.0 565.95 2 3255
Last upload before $\rightarrow$ Event 83.02 20.5 160.84 1 953 129.77 6.0 506.89 1 3187
Event $\rightarrow$ First upload after 72.42 20.5 129.74 1 669 84.49 17.0 172.63 1 903
Figure D.5: Temporal gaps between uploads surrounding reference events. Box plots (top) show the distribution of inactivity intervals for the Attempted (During) and Control (Major Life Event) groups (log scale). The table (bottom) summarizes descriptive statistics for each interval type. The reference event corresponds to a suicide attempt in the Attempted group and a major life event in the Control group.
Appendix E Appendix E - Topic Information
This section includes two tables with prevalent topics identified in this study. Table LABEL:tab:topic_info describes the topics identified in our bottom-up (§3.1) and hybrid approaches (§3.4). Table E.2 presents topics that are highly correlated and have similar semantic meaning. Some information has been reduced or anonymized.
Table E.1: Overview of topic information, including: Count - the number of transcript segments assigned to each topic, Llama2 Name - the name generated by the Llama-7b model (see Section B.1.2), Representative Transcript Segment - the transcript segment closest to the topic’s center and most representative of it, and Top Words - the 10 words with the highest TF-IDF scores. Red topic names indicate topics that were identified in our bottom-up approach. Blue topics were found in our hybrid approach, and brown indicates topics that were found by both approaches. Corresponding * indicates significance with regards to temporal changes in Attempted (During).
@@ Appendix E Appendix E - Topic Information
1041 I’m going to pop in something really sour into my mouth, or I’m going to smell something that’s going to get my brain out of here. Any tips on, uh, anxiety panic attacks, panic attacks are hard to ground yourself from and they’re hard to stop yourself. There is huge differences between anxiety attacks and panic attacks. One of the biggest ones is there’s no trigger necessarily behind a panic attack. Anxiety and Panic Attacks anxiety,panic,attacks,panic attacks,panic attack,attack,anxiety attack,anxious,anxiety attacks,anxiety like
@@
3363 So, how are you today? I’m doing okay. So I can tell you a little bit about Obsessive-Compulsive Disorder, or OCD. It’s a mental illness and it’s one of the anxiety disorders. And so there’s two parts. There’s the obsessions and there’s the compulsions. So the obsessions, those are repetitive and intrusive thoughts. It can be urges or images that cause a lot of anxiety in an individual. And then the second part are the compulsions. Obsessive-Compulsive Disorder (OCD) ocd, compulsions, intrusive, intrusive thoughts, people ocd, ocd like, obsessions, ocd just, obsessive compulsive, obsessive
1036 Hi guys welcome to … I’m recovering from an eating disorder… I’m recovering from an eating disorder. I’ve been in this chair for like two. Yeah well it won’t be two hours after editing. Eating Disorder Recovery [this topic contains private names]
219 She becomes a rebellious sure goes off to college and Cal basically just goes to high school and does his own thing .. It’s so weird … I would say traumatized a little bit Loss and Trauma in a Small Community [this topic contains private names]
@@
Table E.2: Correlation Matrix of 30 Topics Based on Cosine Similarity: The matrix, computed using cosine similarities of the topic embeddings, demonstrates that topics with high correlation are also semantically similar, as anticipated.